OpenStudy (anonymous):

Find all zeros. f(x) = x( x + 2 )( x - 2)( 3^2 - 4)

Directrix (directrix):

>f(x) = x( x + 2 )( x - 2)( 3^2 - 4) Is the last factor supposed to be ( x^2 - 4) rather than ( 3^2 - 4) as shown in the problem?

OpenStudy (anonymous):

the last factor is 3 to the second power subtract 4.

Directrix (directrix):

Zeros are x-intercepts. To get on the x-axis, y must be zero. That means f(x) = 0 0 = x( x + 2 )( x - 2)( 3^2 - 4) By the Zero Product Property, if you get a product of zero, then one or more of the factors is equal to zero.

Directrix (directrix):

0 = x( x + 2 )( x - 2)( 3^2 - 4) 0 = x OR (x + 2 ) = 0 Or (x - 2) = 0 Or ( 3^2 - 4) = 0 Solve each of those for x to get the zeros.

OpenStudy (anonymous):

thank you very much!

Directrix (directrix):

You are welcome. What did you get for the answers?

OpenStudy (anonymous):

i got 3 possible zeros. -2 , 2, 4^2 / 3^2 or 2 and the square root of three over 9... i think that's right, please check and if it's wrong please correct it.

Directrix (directrix):

3 zeros, yes. Get them from the equations. I got x = 0, x = - 2, and x = 2. That's all.

Directrix (directrix):

4^2 / 3^2 or 2 and the square root of three over 9 is a mystery answer. ( 3^2 - 4) = 0 leads nowhere because ( 9 - 4) = 5 which does not equal 0. That is why I asked you early on if that 3^2 was supposed to be x^2. You said no. So, it's a constant term and never equal to 0 in this problem.

OpenStudy (anonymous):

oh okay i see what i did wrong! thank you very much! have a great day or night ^_^

Directrix (directrix):

You are welcome.